OpenStudy (anonymous):

Which is an example of the associative property? A. (x + 3) + 7 = x + (3 + 7) B. x + 5 = 5 + x C. 3 • (x + 5) = 3 • x + 15 D. If x = a, and a = 7, then x = 7.

OpenStudy (shamil98):

I'll give you a hint, the order in which you add or multiply the terms does not matter.

OpenStudy (charlotte123):

https://www.google.com/search?q=associative+property&source=lnms&tbm=isch&sa=X&ei=19i0UtyLJLSA2QWwp4GIDw&ved=0CAcQ_AUoAQ&biw=1366&bih=666 <--- Associative Property Now - Looking at all of this Information - Which one do you think Is the answer?

OpenStudy (anonymous):

I just wanted to know if it was B.....Is it?

OpenStudy (shamil98):

nope.

OpenStudy (charlotte123):

@Diamond_Swag11 Try something with Parenthesis :) Hint hint

OpenStudy (shamil98):

b is the commutative property i meant

OpenStudy (anonymous):

A. Right?

OpenStudy (shamil98):

yes

OpenStudy (shamil98):

(a + b) + c = a + (b + c)

OpenStudy (charlotte123):

YES!

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (anonymous):

and thank you for helping me too

OpenStudy (charlotte123):

Thanks for trying Instead of asking for the answer right away LOL

OpenStudy (shamil98):

bonus round: evaluate: \[\huge \int\limits_{}^{} x^2 e^{x^2}\]

OpenStudy (charlotte123):

No problem!

OpenStudy (shamil98):

i mean integrate*

OpenStudy (charlotte123):

@shamil98 The answer Is G! IT IS ALWAYS G!

OpenStudy (shamil98):

shhh

OpenStudy (charlotte123):

Oops

OpenStudy (shamil98):

noob

OpenStudy (charlotte123):

XD